# The $\theta$-vacuum from functional renormalisation

### v1

@@ Abstract
We study topological properties of a quantum mechanical system with $U(1)$-symmetry within the functional renormalisation group (fRG) approach. These properties include the vacuum energy structure and the topological susceptibility. Our approach works with a complexification of the flow equation, and specifically we embed the original symmetry into the complex plane, $U(1)\rightarrow\mathbb{C}$. We compute the effective potential of a given topological sector by restricting ourselves to field configurations with a given generalised non-trivial Chern–Simons numbers. The full potential is directly constructed from these sector potentials. Our results compare well with the benchmark results obtained from solving the corresponding Schrödinger equation.
@@ I Introduction
We confirm the practical feasibility of our approach by computing the vacuum structure of the model and in particular the vacuum energy and topological susceptibility. Our results, obtained in a local potential approximation, are in good agreement with the benchmark result by numerically solving the vacuum energy directly from the Schrödinger equation. This is illustrated in Figure 1, where we show as the effective potential of the semi-topological model obtained from the fRG approach set up here. This results illustrates the capacity of the fRG approach to accommodate the effects of topological configurations: The effective potential exhibits a non-analytic structure that originates in its semi-topological property, and hence the different topological sectors are clearly visible.
@@ II $U(1)$ symmetric quantum mechanics
$S=\int_{\tau}\left[\frac{m}{2}(\dot{r}^{2}+r^{2}\dot{\vartheta}^{2})-i\frac{\theta}{2\pi}r^{2}\dot{\vartheta}+\frac{g}{4}(r^{2}-1)^{2}\right].$ (4)
@@ II.1 Radial mean field limit: $g\to\infty$
Then, the second term in 1, the ‘$\theta$-term’ becomes a topological term with
@@
$E_{n}(g\to\infty,\theta)=\frac{1}{2m}\left(n-\frac{\theta}{2\pi}\right)^{2}\,.$ (10)
Here, $n\in\mathbb{Z}$ is the quantum number labelling the energy spectrum, arising from the quantisation of angular fluctuations. For its derivation, see Section A.1. The ground-state energy in this system is periodic in $\theta$, as shown in the panel on the left side of Figure 2. This fact indicates the topological nature of this model. We also note for later use that the normalisation of the radial part to $r=1$ is a pure convention; the above derivation holds for any radial mean field value $r=\bar{r}$ instead of $r=1$. In the more general case, the energy levels 10 turn into
$E_{n}(g\to\infty,\theta)=\frac{1}{2m}\left(n-\frac{\theta\,\bar{r}^{2}}{2\pi}\right)^{2}\,.$ (11)
@@ II.2 General case with finite couplings $g$
$\psi(r,\vartheta)=\psi(r)\psi(\vartheta)\,.$ (12)
The angular variable $\vartheta$ does not enter the $\theta$-term and the potential. Hence, the solution of the Schrödinger equation for the wave function $\psi(\vartheta)$ is given by $\psi_{n}(\vartheta)=e^{in\vartheta}$ up to an overall factor. The Schrödinger equation for $\psi(r)$ reads
@@
$\hskip 28.45274pt+\frac{g}{4}(r^{2}-1)^{2}\Biggr{]}\psi_{l,n}(r)=E_{l,n}(g,\theta)\psi_{l,n}(r)\,.$ (13)
see also Section A.2. If solving this equation with the boundary condition $\psi_{l,n}(r\to\infty)\to 0$, we obtain the energy eigenvalues $E_{l,n}(g,\theta)$. Here, $l$ stands for the quantum number corresponding to the radial excitation. In this work, we consider the ground state $l=0$ and denote
$E_{n}(g,\theta):=E_{l=0,n}(g,\theta)\,.$ (14)
@@
$E_{n}(g=0,\theta)=\frac{|\theta|}{\pi m}\left(n+\frac{1}{2}\right)\,.$ (15)
@@ III Functional renormalisation group approach to $U(1)$-symmetric quantum mechanics
In the present work we mostly consider the Cartesian field $\hat{\phi}=\hat{\varphi}$, specifically within the numerical applications. However, for the discussion of potential regularisation of the topological degrees of freedom we also consider the polar coordinates $\hat{\phi}[\hat{\varphi}]=(\hat{r},\hat{\vartheta})$.
@@ III.2.2 Initial effective action
×$\hskip 71.13188pt\times e^{\int_{\tau}\left(\hat{\varphi}-\varphi\right)\frac{\delta\Gamma[\varphi]}{\delta\varphi}}\,.$ (27)
Assume now that $\varphi$ is a fluctuation mean field in the trivial topology sector. Then, even though $[\mathcal{D}\hat{\varphi}]$ integrated over all fields, it is reduced to the measure in the trivial topology sector due to the regulator term: for $\hat{\varphi}$ with non-trivial windings the regulator term diverges in the limit $\beta\to\infty$.
@@
We show the potential 32b as a function of $\theta$ for $r^{2}=1$ for different inverse temperatures $\beta$. In particular, in the limit $\beta\to\infty$ we obtain
@@
$=$ $\frac{1}{2m}\left(\frac{\theta}{2\pi}r-\frac{n(\theta,r)}{r}\right)^{2}\,,$ (33)
@@
The potential 33 in the $r^{2}$-$\theta$ plane is shown in Figure 4, for the slice $V_{\textrm{top},\beta}(r^{2},\theta=2\pi)$ see Figure 5. Equation 33 entails that this potential has several non-analytical cusps that separate the $r^{2}$-$\theta$-plane into different branches. These cups originate from the transition between different “energy levels” 34 when taking the limit $\beta\rightarrow\infty$. There, they can be read read off as
@@ III.3 Flow equation of the effective potential
Using 40, we arrive at the following analytic flow equation for the effective potential, see 82.
@@ IV.1 Energy levels
In 44 we have dropped an additive constant linear in $k$. This term takes into account that the overall normalisation of the path integral is $k$-dependent and has to chosen such that at $k=0$ the states are normalised to unity: 1=1$\langle$ 1$\rangle$=1⟨ 1 ⟩ = 1. Then, the effective potential will not only allow us to compute energy difference but also the vacuum energy. The constant term does not feed back into the flow and we simply remove the respective contributions in the flow,
@@
Evidently, the subtraction in 45 is independent of the potential $V_{k}$ and only removes the cutoff dependence of the constant term. With a slight abuse of notation we call the subtracted potential still $V_{k}$.
@@
$E_{n}(\theta;g)=V_{k=0}(r^{2}=\langle r\rangle^{2})\,,$ (46)
where $\langle r\rangle$ denotes the minimum of the effective potential. Finally, we note that in quantum mechanical systems (one-dimensional quantum field theories), we always observe $\langle r\rangle=0$ thanks to the Hohenberg-Mermin-Wagner theorem Mermin:1966fe ; Hohenberg:1967zz , see also Berezinskii and Coleman Berezinsky:1970fr ; Coleman:1973ci .
Figure 6: Energy eigenvalues (dots) for $n=0$ in terms of the $\theta$-parameter from the flow equations with with the Litim-type regulator. The solid lines show the solution of the Schrödinger equation 13.
@@
[…]
@@
Figure 7: Energy levels in terms of the $\theta$-parameter from the flow equations with the Litim-type regulator. The solid lines show the solution of the Schrödinger equation 13.
IV.2 Evolution of effective potential with cusps
We proceed with the computation of the full effective potential $V(r^{2},\theta)$ without any topological restrictions. This is done by starting the flow from the initial effective action in 32. For the sake of completeness we remark that this can be done directly for all $r$ and $\theta$ if we would monitor the position of the cusps during the flow: the flow in between the cusps is that of $V_{n}$ for a given $n$. In Appendix F we show how this is done practically; see in particular 102 and 99.
@@ IV.2 Evolution of effective potential with cusps
However, while possible, the extraction of $V_{k}(r^{2},\theta)$, does not necessitate new numerical computations, but simply the use of the effective potentials $V_{n}$ obtained in the different topological sectors: as mentioned above, within the cusps the flow of $V_{k}$ is simply that of $V_{n,k}$ for a given $n$. The flowing cusps are simply the intersection points (or level crossings) of the potentials $V_{n}$. Hence, the complete potential $V(r^{2},\theta)$ is obtained by the following definition
$V(r^{2},\theta)=\min_{n}V_{n}(r^{2},\theta)\,.$ (47)
@@
Figure 9: Effective potential $V(r^{2},\theta=2\pi)$ evaluated at $k=3000$ (dashed) and $k=0.001$ (solid), from the flow equation with Litim-type regulator 42. Different colors of the potential characterise the different branches of the quantum number $n$ defined in 44.
@@ IV.3 Topological susceptibility
The results obtained from the Schrödinger equation are shown in Figure 10. For large $g\geq 1$, $I_{\rm top}(g)$ approaches 49. Moreover, it features a minimum at $g=11.75$. In Figure 10 we compare the exact results with those obtained from the flow equation with the Litim regulator 42.
@@
We close this Section with the remark that an obvious and interesting extension of the present analysis is the computation $\langle\dot{\varphi}(t)\dot{\varphi}(0)\rangle$. For this computation, one has to go beyond the LPA-approximation used in the present work, and respective results will be presented elsewhere.
@@ V Summary and Conclusion
Figure 10: Topological susceptibility 50 with respect to the constraint coupling $g$. The LPA-fRG result (circles) is obtained by the Litim-type flow equation 42. The red line is the result from the Schrödinger equation 13. The dashed line exhibits the value 49 in topological quantum mechanics.
@@
In summary, our work demonstrates the potential of fRG to address topological aspects of quantum systems in a well-controlled but non-trivial benchmark theory. Our results complements those in the anharmonic oscillator in Bonanno:2025mon . The analysis in the present work also provides a road map for the application of our findings and techniques to the $\theta$-vacuum structure in gauge theories. While we specifically aim at QCD and respective computations are under way, a first step in this programme is provided by the study of the 2D Schwinger model. We hope to report on results for both cases soon.
@@ A.1 Quantum rotor and its topological vacuum structure
where the dot represents the derivative with respect to $\tau$. The field variable $z=e^{i\vartheta}$ is an element of $U(1)$. As shown in 6, the second term in 51 corresponds to a topological term. The path integral for this system is given by
@@ A.2 Relaxation to a $U(1)$-symmetric theory
see also 32 with $\hat{\phi}=\hat{\varphi}$. For finite coupling $g$, the quantum rotor with the generating functional $Z_{z}$ in 52 is deformed into a $U(1)$-symmetric model with the complex field $\varphi$. In terms of the fields $r$ and $\vartheta$, the classical action 60 is given by
@@
The respective Hamiltonian follows with a Legendre transformation from the Lagrangian in 63 as
@@ Appendix B Derivation of the flow equation in the local potential approximation
[…]
@@
The flow equation of the effective potential with the Litim-type regulator follows from 80 and 81 as
@@ Appendix C Reference computations with the CS-regulator
[…]
@@
Figure 11: Energy eigenvalues $E_{n}$ (dots) as functions of the $\theta$-parameter obtained from the flow equations with the CS regulator 86. The left panel shows $E_{0}$ for various values of $g$, while the right panel displays $E_{n}$ for several values of $n$ at fixed $g=30$. The solid lines show the solution of the Schrödinger equation 13.

### v2

@@ Abstract
We study topological properties of a quantum mechanical system with $U(1)$-symmetry within the functional renormalisation group (fRG) approach. These properties include the vacuum energy structure and the topological susceptibility. Our approach works with a complexification of the flow equation, and specifically we embed the original symmetry into the complex plane, $U(1)\rightarrow\mathbb{C}$. We compute the effective potential of a given topological sector by restricting ourselves to field configurations with given generalised non-trivial Chern–Simons numbers. The full potential is directly constructed from these sector potentials. Our results compare well with the benchmark results obtained from solving the corresponding Schrödinger equation.
@@ I Introduction
We confirm the practical feasibility of our approach by computing the vacuum structure of the model and, in particular, the vacuum energy and topological susceptibility. Our results, obtained in a local potential approximation, are in good agreement with the benchmark result by numerically solving the vacuum energy directly from the Schrödinger equation. This is illustrated in Figure 1, where we show as the effective potential of the semi-topological model obtained from the fRG approach set up here. This result illustrates the capacity of the fRG approach to accommodate the effects of topological configurations: The effective potential exhibits a non-analytic structure that originates in its semi-topological property, and hence the different topological sectors are clearly visible.
@@ II $U(1)$ symmetric quantum mechanics
$S=\int_{\tau}\left[\frac{m}{2}(\dot{r}^{2}+r^{2}\dot{\vartheta}^{2})-i\frac{\theta}{2\pi}r^{2}\dot{\vartheta}+\frac{g}{4}(r^{2}-1)^{2}\right],$ (4)
@@ II.1 Radial mean field limit: $g\to\infty$
Then, the second term in 1, the $\theta$-term’ becomes a topological term with
@@
$E_{n}(\theta;g\to\infty)=\frac{1}{2m}\left(n-\frac{\theta}{2\pi}\right)^{2}\,.$ (10)
Here, $n\in\mathbb{Z}$ is the quantum number labeling the energy spectrum, arising from the quantisation of angular fluctuations. For its derivation, see Section A.1. The ground-state energy in this system is periodic in $\theta$, as shown in the panel on the left side of Figure 2. This fact indicates the topological nature of this model. We also note for later use that the normalisation of the radial part to $r=1$ is a pure convention; the above derivation holds for any radial mean field value $r=\bar{r}$ instead of $r=1$. In the more general case, the energy levels 10 turn into
$E_{n}(\theta;g\to\infty)=\frac{1}{2m}\left(n-\frac{\theta\,\bar{r}^{2}}{2\pi}\right)^{2}\,.$ (11)
@@ II.2 General case with finite couplings $g$
$\psi(r,\vartheta)=\sum_{l,n}\Upsilon_{l,n}(r)\Phi_{n}(\vartheta)\,.$ (12)
The angular variable $\vartheta$ does not enter the $\theta$-term and the potential. Hence, the solution of the Schrödinger equation for the wave function $\Phi_{n}(\vartheta)$ is given by $\Phi_{n}(\vartheta)=e^{in\vartheta}$ up to an overall factor. The Schrödinger equation for $\Upsilon_{l,n}(r)$ reads
@@
$\hskip 28.45274pt+\frac{g}{4}(r^{2}-1)^{2}\Biggr]\Upsilon_{l,n}(r)=E_{l,n}(\theta;g)\Upsilon_{l,n}(r)\,$, (13)
see also Section A.2. If solving this equation with the boundary condition $\Upsilon_{l,n}(r\to\infty)\to 0$, we obtain the energy eigenvalues $E_{l,n}(\theta;g)$. Here, $l$ stands for the quantum number corresponding to the radial excitation. In this work, we consider the ground state $l=0$ and denote
$E_{n}(\theta;g):=E_{l=0,n}(\theta;g)\,.$ (14)
@@
$E_{n}(\theta;g=0)=\frac{|\theta|}{\pi m}\left(n+\frac{1}{2}\right)\,.$ (15)
@@ III Functional renormalisation group approach to $U(1)$-symmetric quantum mechanics
[…]
@@ III.2.2 Initial effective action
$\hskip 71.13188pt\times e^{\int_{\tau}\left(\hat{\varphi}-\varphi\right)\frac{\delta\Gamma[\varphi]}{\delta\varphi}}\,.$ (27)
Assume now that $\varphi$ is a fluctuation mean field in the trivial topology sector. Then, even though $\mathcal{D}\hat{\varphi}$] integrated over all fields, it is reduced to the measure in the trivial topology sector due to the regulator term: for $\hat{\varphi}$ with non-trivial windings the regulator term diverges in the limit $\beta\to\infty$.
@@
We show the potential 32b as a function of $\theta$ for $r^{2}=1$ for different inverse temperatures $\beta$ in Figure 3. In particular, in the limit $\beta\to\infty$ we obtain
@@
$=$ $\frac{1}{2m}\left(\frac{\theta}{2\pi}r^{2}-\frac{n(\theta,r)}{r}\right)^{2}\,$, (33)
@@
The potential 33 in the $r^{2}$-$\theta$ plane is shown in Figure 4, for the slice $V_{\textrm{top},\beta}(r^{2},\theta=2\pi)$ see Figure 5. Equation 33 entails that this potential has several non-analytical cusps that separate the $r^{2}$-$\theta$-plane into different branches. These cusps originate from the transition between different “energy levels” 34 when taking the limit $\beta\rightarrow\infty$. There, they can be read off as
@@ III.3 Flow equation of the effective potential
Using 40, we arrive at the following analytic flow equation for the effective potential, see 83.
@@ IV.1 Energy levels
In 44 we have dropped an additive constant linear in $k$. This term takes into account that the overall normalisation of the path integral is $k$-dependent and has to be chosen such that at $k=0$ the states are normalised to unity: $\langle 1\rangle=1$. Then, the effective potential will not only allow us to compute energy difference but also the vacuum energy. The constant term does not feed back into the flow and we simply remove the respective contributions in the flow,
@@
Evidently, the subtraction in 45 is independent of the potential $V_{k}$, and hence on the energy level. With a slight abuse of notation we still call the subtracted potential $V_{k}$.
The subtraction in 45 removes the leading cutoff dependence of the field-independent part of the potentials $V_{n,k}$. The regulator also induces further ones which lower powers of $k$. Note however that they may also feature a $\theta$- and $n$-dependence which is absent in the leading order term. This will be discussed elsewhere. In the current work we circumvent this issue as follows: we normalise the fRG results for the energy levels $E_{n}(\theta;g)$ to that from the Schrödinger equations at the minimum $\theta_{\textrm{min},n}$.
@@
$E^{\textrm{fRG}}_{n}(\theta;g)=$ $\,V_{n,k=0}(\langle r\rangle^{2},\theta)-V_{n,k=0}(\langle r\rangle^{2},\theta_{\textrm{min},n})$
$+E^{\textrm{Ham}}_{n}(\theta_{\textrm{min},n};g)\,.$ (46)
Here, $\langle r\rangle$ denotes the minimum of the effective potential and $E^{\textrm{fRG/Ham}}$ denotes the energy levels computed with the fRG and the Schrödinger equation respectively. Finally, we note that in quantum mechanical systems (one-dimensional quantum field theories), we always observe $\langle r\rangle=0$ thanks to the Hohenberg-Mermin-Wagner theorem Mermin:1966fe ; Hohenberg:1967zz , see also Berezinskii and Coleman Berezinsky:1970fr ; Coleman:1973ci . eq. 46
Figure 6: Energy levels 46 for $n=0$ and different couplings $g$ in terms of the $\theta$-parameter from the flow equations with the Litim-type regulator. The solid lines show the energy levels obtained from the Schrödinger equation 13.
@@
For a fixed value $g=30$, each energy level (n=0, 1, 2, 3, 4n=0,\,1,\,2,\,3,\,4) is displayed as a function of $\theta$ in Figure 7. One finds that the $\theta$-dependence is well captured for each quantum number $n$, or equivalently the winding number $\nu$. However, in this approach, the level crossing in the ground-state energy is not observed because of the fixed topology (quantum number).
@@
Figure 7: Energy levels 46 for different $n$’s and $g=30$ in terms of the $\theta$-parameter from the flow equations with the Litim-type regulator. The solid lines show the energy levels obtained from the Schrödinger equation 13.
IV.2 Evolution of the effective potential with cusps
We proceed with the computation of the full effective potential $V(r^{2},\theta)$ without any topological restrictions. This is done by starting the flow from the initial effective action in 32. For the sake of completeness we remark that this can be done directly for all $r$ and $\theta$ if we would monitor the position of the cusps during the flow: the flow in between the cusps is that of $V_{n}$ for a given $n$. In Appendix F we show how this is done practically; see in particular 106 and 103.
@@ IV.2 Evolution of the effective potential with cusps
However, while possible, the extraction of $V_{k}(r^{2},\theta)$, does not necessitate new numerical computations, but simply the use of the effective potentials $V_{n}$ obtained in the different topological sectors: as mentioned above, within the cusps the flow of $V_{k}$ is simply that of $V_{n,k}$ for a given $n$. The flowing cusps are simply the intersection points (or level crossings) of the potentials $V_{n}$.
In the present work we are only interested in the potential at $k=0$. There we exploit Equation 46 which already includes the relative normalisation between the sectors. Then, the complete potential $V(r^{2},\theta)$ is obtained by the following definition
$V(r^{2},\theta)=$ $\,\min_{n}\Bigl[V_{n,k=0}(r^{2},\theta)-V_{n,k=0}(\langle r\rangle^{2},\theta)$
$\hskip 28.45274pt+E_{n}^{\textrm{fRG}}(\theta;g)\Bigr]\,.$ (47)
@@
Equation 47 leads us to the final expression for the energy of the system,
$E(\theta;g)=V(r^{2}=\langle r\rangle^{2})=\min_{n}E_{n}^{\textrm{fRG}}(\theta;g)\,.$ (49)
Within the current approximation, this result is displayed in Figure 10 together with the exact numerical result.
Figure 9: Effective potential $V(r^{2},\theta=2\pi)$ evaluated at $k=3000$ (dashed) and $k=0.001$ (solid), from the flow equation with Litim-type regulator 42. Different colors of the potential characterise the different branches of the quantum number $n$ defined in $4$.
@@ IV.3 Topological susceptibility
The results obtained from the Schrödinger equation are shown in Figure 11. For large $g\geq 1$, $I_{\rm top}(g)$ approaches 50. Moreover, it features a minimum at $g=11.75$. In Figure 11 we compare the exact results with those obtained from the flow equation with the Litim regulator 42.
@@
We close this Section with the remark that an obvious and interesting extension of the present analysis is the computation $\langle\dot{\varphi}(t)\dot{\varphi}(0)\rangle$. For this computation, one has to go beyond the LPA-approximation used in the present work, and the respective results will be presented elsewhere.
Figure 10: Ground-state energy 49. We display results obtained with the flow equation with the CS-type regulator (dashed blue) and with the Litim-type regulator (dashed red). The solid gray line denotes the benchmark result obtained by solving the eigenvalue problem of the Schrödinger equation 13. The colored dotted lines indicate the energies of the excited states, as shown in the right panel of Figure 2.
@@ V Summary and Conclusion
Figure 11: Topological susceptibility 51 with respect to the constraint coupling $g$. The LPA-fRG result (circles) is obtained by the Litim-type flow equation 42. The red line is the result from the Schrödinger equation 13. The dashed line exhibits the value 50 in topological quantum mechanics.
@@
In summary, our work demonstrates the potential of fRG to address topological aspects of quantum systems in a well-controlled but non-trivial benchmark theory. Our results complement those in the anharmonic oscillator in Bonanno:2025mon . The analysis in the present work also provides a road map for the application of our findings and techniques to the $\theta$-vacuum structure in gauge theories. While we specifically aim at QCD and respective computations are under way, a first step in this programme is provided by the study of the 2D Schwinger model. We hope to report on results for both cases soon.
@@ A.1 Quantum rotor and its topological vacuum structure
where the dot represents the derivative with respect to $\tau$. The field variable $z=e^{i\vartheta}$ is an element of $U(1)$. As shown in 6, the second term in 52 corresponds to a topological term. The path integral for this system is given by
@@ A.2 Relaxation to a $U(1)$-symmetric theory
see also 32 with $\hat{\phi}=\hat{\varphi}$. For finite coupling $g$, the quantum rotor with the generating functional $Z_{z}$ in 53 is deformed into a $U(1)$-symmetric model with the complex field $\varphi$. In terms of the fields $r$ and $\vartheta$, the classical action 61 is given by
@@
The respective Hamiltonian follows with a Legendre transformation from the Lagrangian in 64 as
@@ Appendix B Derivation of the flow equation in the local potential approximation
$\hskip-56.9055pt=\frac{1}{2}{\rm Tr}\,\left[\left(\begin{array}[]{cc}\varphi^{*2}V_{k}^{\prime\prime}&\frac{m}{2}P_{k}(p)+V_{k}^{\prime}+r^{2}V_{k}^{\prime\prime}\\[8.61108pt] \frac{m}{2}P_{k}(p)+V_{k}^{\prime}+r^{2}V_{k}^{\prime\prime}&\varphi^{2}V_{k}^{\prime\prime}\end{array}\right)^{-1}\begin{pmatrix}0&\frac{m}{2}\partial_{t}{R}_{k}(p)\\[4.30554pt] \frac{m}{2}\partial_{t}{R}_{k}(p)&0\end{pmatrix}\,2\pi\,\delta(p+q)\right]$ (78)
@@
The flow equation of the effective potential with the Litim-type regulator follows from 81 and 82 as
@@ Appendix C Reference computations with the CS-regulator
We have solved the flow equation with the CS-regulator 87 and computed the energy levels 46. The left-hand side panel of Figure 12 depicts the ground-state energy $E_{0}(\theta;g)$ with several values of $g$ as a function of $\theta$. In the right-hand side panel of Figure 12, we also show the $\theta$-dependence of the energy levels (n=0, 1, 2, 3, 4n=0,\ 1,\ 2,\ 3,\ 4) with $g=30$.
@@
Figure 12: Energy eigenvalues $E_{n}$ (markers) as functions of the $\theta$-parameter obtained from the flow equations with the CS regulator 87. The left panel shows $E_{0}$ for various values of $g$, while the right panel displays $E_{n}$ for several values of $n$ at fixed $g=30$. The solid lines show the solution of the Schrödinger equation 13.